# Deep Neural Networks: Multi-Classification and Universal Approximation

### v1

@@ 5. Universal approximation theorem
In this section, we prove the Universal Approximation Theorem in $L^{p}(\Omega;\mathbb{R}_{+})$.
Proof of Theorem 1.2..
@@
with $C_{\Omega}$ a constant depending on $m_{d}(\Omega)$, where $m_{d}(\cdot)$ is the Lebesgue measure in $\mathbb{R}^{d}$. Taking into account that the number of edges of a $d$-dimensional hypercube is $2d(d-1)$, the Lebesgue measure $m_{d}(G^{h}_{\delta})$ of the grid $G_{\delta}^{h}$ intersecting $\mathcal{C}$ is bounded by
@@
Thus, for any $\gamma>0$, taking $\delta=h^{1+\gamma}$, the volume of $G_{\delta}^{h}$ tends to zero as $h\rightarrow 0$.
Figure 20. (A) Illustration of the level sets of the function $f(x,y)=x^{2}+y^{2}$ defined on $\Omega\subset\mathbb{R}^{2}$. The rectangle represents the set $\mathcal{C}$. (B) We represent the function $f$ extended by zero to $\mathcal{C}$. (C) Representation of the level set of $f$ on $\mathcal{H}=\mathcal{C}\setminus G_{\delta}^{h}$. Here the white mesh represent $G_{\delta}^{h}$
Figure 21. (A) Illustration of the simple function $f_{h}$ on $\mathcal{H}$, defined in (5.3). (B) The main features of $G^{h}_{\delta}$ are represented: The mesh thickness $\delta$, the size $h>0$, and the hyperrectangles $\mathcal{H}^{G}_{i}$.
@@
That is, $f_{i}^{h}$ is the average value of the function $f$ in the hyperrectangle $\mathcal{H}_{i}$. Then, we introduce the simple function
$f_{h}(x)=\sum_{i=1}^{N_{h}}f_{i}^{h}\chi_{\mathcal{H}_{i}}(x)+\sum_{i=1}^{N_{h}^{G}}f_{i}^{h}\chi_{\mathcal{H}_{i}^{G}}(x),$ (5.3)
where $\chi_{\mathcal{H}_{i}}$ denotes the characteristic function on the set $\mathcal{H}_{i}$. In the following, we denote by $M_{h}>0$ the number of values that the function $f_{h}$ takes on the family of hyperrectangles $\mathcal{H}$. Note that $M_{h}\leq N_{h}$.
@@
Moreover, as shown in [14, Section 6.2], if $f\in W^{1,p}(\Omega;\mathbb{R}_{+})$, there exists a constant $C>0$ independent of $h>0$ such that
@@
where $\text{diam}(\mathcal{H})=\max_{\mathcal{H}_{i}\in\mathcal{H}}\{\text{diam}(\mathcal{H}_{i})\}$. Since $\text{diam}(\mathcal{H^{G}})<\text{diam}(\mathcal{H})$, and $\text{diam}(\mathcal{H})$ is at most $h>0$, (5.5) reduces to
@@
$h_{1}\leq\frac{\varepsilon}{2C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}.$ (5.7)
@@
Here, we have one hyperplane for each canonical direction and one extra hyperplane in the direction $e_{\hat{\eta}}$, so there are in total $d+1$ hyperplanes. The constants $b_{k}$ are chosen in such a way that the hyperplanes are placed around the hyperrectangle $\mathcal{H}_{*}$, that is, $b_{k}$ is taken such that
$\begin{cases}\sigma(e_{\eta}x+b_{\eta})=0,\\ \sigma(-e_{\hat{\eta}}x+b_{\hat{\eta}}^{*})=0,\end{cases}\quad\text{for all }x\in\mathcal{H}_{*},\,\eta\in\llbracket 1,2d\rrbracket.$
Let us apply the hyperplanes (5) to $\mathcal{H}$. This defines the new family of hyperrectangles $\{\mathcal{H}_{i}^{1}\}_{i=1}^{N}$. Namely, denote by $W^{1}\in\mathbb{R}^{d+1\times d}$ and $b^{1}\in\mathbb{R}^{d+1}$ the matrices given by
@@
Since (5.10), we have that
$p_{\hat{\eta}}=0\quad\text{or}\quad p_{d+1}=0.$ (5.13)
@@
We have that $F(\mathcal{H}_{*})={\bf 0}_{d}$, where ${\bf 0}_{d}$ denotes the null vector in $\mathbb{R}^{d}$. Moreover, for any hyperrectangles $\mathcal{H}_{j},\mathcal{H}_{i}\subset\mathcal{H}$ with $i\neq j$, we have that $F(\mathcal{H}_{i})\cap F(\mathcal{H}_{j})=\emptyset$.
@@
$=\left(\sum_{\eta=1}^{d}p_{\eta}(x)e_{\eta}x+p_{d+1}(x)e_{\hat{\eta}}x\right)+\left(\sum_{\eta=1}^{d}p_{\eta}(x)b_{\eta}e_{\eta}+p_{d+1}(x)b_{\hat{\eta}}e_{\hat{\eta}}\right)+b^{2}=F(x)+b^{2}.$
@@
Let $E$ represent the set of edges of $\mathcal{C}$. We define $\{E_{i}\}_{i=1}^{d}\subset E$ as a set of orthogonal edges. The family of hyperrectangles $\mathcal{H}^{E}$ is then given by $\mathcal{H}^{E}=\{\mathcal{H}_{i}\in\mathcal{H}\,:\,\mathcal{H}_{i}\cap\{E_{i}\}_{i=1}^{d}\neq\emptyset\},$ that is, the set of hyperrectangles in $\mathcal{H}$ that intersect at least one edge in $\{E_{i}\}_{i=1}^{d}$. We will prove that compressing the hyperrectangles in $\mathcal{H}^{E}$ into points also compresses all the hyperrectangles in $\mathcal{H}$ into points.
Figure 23. In the figure, we show an example of the compression process following the guidelines from Steps 2.1-2.2. First, in Step 1, we select two edges, $E_{1}$ and $E_{2}$ of $\mathcal{C}$, oriented according to different canonical vectors. Thus, $\mathcal{H}^{E}=\{\mathcal{H}_{1},\mathcal{H}_{4},\mathcal{H}_{7},\mathcal{H}_{8},\mathcal{H}_{9}\}$. Then, we compress the first hyperrectangle, $\mathcal{H}1$, using two hyperplanes (a two-layer neural network with two neurons in the first layer and two neurons in the second layer). Here, $l{1,1}^{E_{1}}$ is a hyperplane that intersects $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$, and $l_{2,1}^{E_{1}}$ intersects $\mathcal{H}_{4}$ and $\mathcal{H}_{7}$ (observe that these hyperrectangles were mapped into lines after the second step). Next, in Step 2, we compress the second hyperrectangle on $E_{1}$, corresponding to $\mathcal{H}4^{2}$. We use three hyperplanes to compress it (a two-layer neural network with three neurons in the first layer and two neurons in the second layer). Here, $l_{2,2}^{E_{1}}$ intersects $\mathcal{H}_{5}^{2}$ and $\mathcal{H}_{6}^{2}$. In Step 3, we continue with $\mathcal{H}_{7}^{4}$. After compressing all the hyperrectangles along $E_{1}$, we have reduced every hyperrectangle to objects of $(d-1)$ dimensions ($d-1=1$ in this example). By applying the same process to the hyperrectangles along $E_{2}$ and continuing with Steps 4 and 5, we complete the compression process.
@@
Let $E_{1}$ be an edge of $\mathcal{C}$. Thanks to Step 2.1, we can compress the first hyperrectangle $\mathcal{H}_{1}$ along $E_{1}$ using a two-layer neural network. Moreover, consider a point $c\in\mathcal{H}_{1}$. Due to the construction of the hyperplanes in Step 2.1, all hyperrectangles that intersect the hyperplanes
$l_{j,1}^{E_{1}}:=\{e_{j}x+c=0\}\text{ for }j\in\llbracket 1,d\rrbracket,$
are mapped into hyperrectangles of at most $(d-1)$ dimensions, as these hyperrectangles are confined within at most $d-1$ regions. Here, $l_{j,k}^{E_{1}}$ denotes a hyperplane in the canonical $j$-th direction, where $c$ belongs to the $k$-th hyperrectangle over $E_{1}$.
After applying Step 2.1 to all hyperrectangles on edge $E_{1}$, all hyperrectangles ${\mathcal{H}i}{i=1}^{N_{h}}$ that were intercepted by at least one $l_{j,k}^{E_{1}}$ for some $j$ and $k$ are mapped to $(d-1)$-dimensional hyperrectangles.
Next, we proceed with the hyperrectangles on $E_{2}$. In this step, the dimension of all hyperrectangles in a subregion intersecting any $l_{j,k}^{E_{2}}$ is further reduced by one. Specifically, if a $(d-1)$-dimensional hyperrectangle was in a subregion where an $l_{j,k}^{E_{2}}$ passed, it is now mapped to a $(d-2)$-dimensional hyperrectangle.
By applying this process inductively to all hyperrectangles in $\mathcal{H}^{E}$, all the hyperrectangles in $\mathcal{H}$ are eventually reduced to 00-dimensional hyperrectangles, i.e., points. This concludes the compression process.
@@
Let us observe that the previous step is essential for achieving the optimal result. This step allows us to reduce not only the number of steps but also the number of hyperplanes used in each step. Note that it was only necessary to compress the hyperrectangles on the edges of $\mathcal{C}$, and for this, we used $d+1$ hyperplanes at each step. However, if we needed to compress a hyperrectangle inside $\mathcal{C}$, we would require $2d$ hyperplanes, which would increase the width of the neural network.
@@
$\phi^{2N_{E}}(\mathcal{W}^{2N_{E}},\mathcal{B}^{2N_{E}},\mathcal{H})=\{x_{i}\}_{i=1}^{N_{h}},$ (5.17)
@@
$\phi^{L}(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=f^{h}_{i},\quad\text{for all }i\in\llbracket 1,N_{h}\rrbracket.$ (5.18)
@@
$\phi^{\mathcal{L}}(\mathcal{W}^{\mathcal{L}},\mathcal{B}^{\mathcal{L}},x)=f_{h}(x),\quad\text{for every }x\in\mathcal{H},$ (5.19)
@@
$\leq 2^{p}C_{1}^{p}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})\sum_{i=1}^{N_{h}^{G}}m_{d}(\mathcal{H}_{i}^{G})=2^{p}C_{1}^{p}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})m_{d}(G^{h}_{\delta}).$
@@
with $C_{2}=2^{p}C_{1}^{p}C_{\Omega,d}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})$. Then, taking into account that $\delta<h^{1+\gamma}$, for any $\gamma>0$, we obtain
@@
Let $\phi^{\mathcal{L}}$ be the map defined by (5.19), and denote by $l_{\mathcal{C}}$ the longest edge of $\mathcal{C}$. Then, for $h<l_{\mathcal{C}}\log(2)/(dp+1)$ we have $\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}^{p}\leq C_{3}\left(1+\delta^{p}(h+\delta)+h^{p}\right),$ with $C_{3}>0$ a constant that depends on $d$, $m_{d}(\mathcal{C})$, $p$ and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$.
Let us assume that $h<l_{\mathcal{C}}\log(2)/(dp+1)$. Then, using Lemma 5.2, (5.2), and since $\delta<h^{\gamma+1}$ we deduce
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(G_{\delta}^{h};\mathbb{R}_{+})}^{p}m_{d}(G^{h}_{\delta})$ $\leq C_{3}\left(1+\delta^{p}(h+\delta)+h^{p}\right)C_{\Omega,d}\delta(h+\delta)^{d-1}h^{-d}\leq 4C_{3}C_{\Omega,d}\delta.$ (5.23)
@@
$\|\phi^{\mathcal{L}}-f_{h}\|^{p}_{L^{p}(\Omega;\mathbb{R}_{+})}$ $\leq C_{2}\delta+4C_{3}C_{\Omega,d}\delta\leq\hat{C}\delta,$ (5.24)
where $\hat{C}=C_{2}+4C_{3}C_{\Omega,d}$ is a positive constant depending on $d$, $p$, $m_{d}(\mathcal{C})$, and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$. Thus, taking $\delta<\varepsilon^{p}/(\hat{C}2^{p})$, we deduce that
$\|\phi^{\mathcal{L}}-f_{h}\|^{p}_{L^{p}(\Omega;\mathbb{R}_{+})}<\frac{\varepsilon}{2}.$ (5.25)
Therefore, taking $\delta<\min\{\varepsilon^{p}/(\hat{C}2^{p}),h_{1}^{1+\gamma}\}$, we have that (5.4) and (5.24) hold, and consequently, we conclude that $\|f-\phi^{\mathcal{L}}\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon$.
Step 4 (Depth Estimation): To estimate the depth of the neural network, it is necessary to estimate $\delta$ such that $\delta<\min\{\varepsilon^{p}/(\hat{C}2^{p}),h^{\gamma+1}_{1}\}$. Assume that $f\in W^{1,p}(\Omega;\mathbb{R}_{+})$. Then, let us observe that in an analogous way to (5.6), we can deduce that
@@
Therefore, using (5.26) instead (5.21), we can consider $C_{2}=2^{p}C_{1}^{p}C_{\Omega,d}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$. Moreover, due Lemma 5.2, we have that $C_{3}=C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$. Consequently, we deduce that the constant $\hat{C}$ can be taken as $\hat{C}=(2^{p}C_{1}^{p}C_{\Omega,d}+4CC_{\Omega,d})\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$.
On the other hand, to estimate $\delta$, we have seen in Step 1, that if $f\in W^{1,p}(\Omega;\mathbb{R}_{+})$ then $h_{1}\leq\varepsilon/(C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})})$, ensures (5.4). Therefore, considering $\delta\leq h^{1+p}$ with $h<h_{1}<1$, to ensure that $\|f-\phi^{\mathcal{L}}\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon$, we need to take $\delta<\frac{\varepsilon^{p}}{2^{p}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}}\min\{1/\hat{C},\varepsilon/(2C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})})\}$. Then, in particular, we take
$\delta=\frac{C_{5}\varepsilon^{p+1}}{2^{p+1}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p+1}},$ (5.27)
with $C_{5}=\min\{1/\hat{C},1/C\}$. Now, due to Step 2.3, we know that $\mathcal{L}=2N_{h}+4M_{h}-1+2N_{E}$. Moreover, observe that $N_{E}$ and $N$ can be estimated by
@@
Then, using $M_{h}\leq N_{h}$ and the estimations (5.28) and (5.27), we deduce the upper bound for the depth
$\mathcal{L}\leq C_{6}\left(\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{dp}\varepsilon^{-dp}+\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}\varepsilon^{-p}\right)+7,$

### v2

@@ 5. Universal approximation theorem
In this section, we prove the Universal Approximation Theorem in $L^{p}(\Omega;\mathbb{R}_{+})$ (Theorem 1.4).
Proof.
@@
$G_{\delta}^{h}=\bigcup_{i=1}^{N_{G}}\mathcal{H}_{i}^{G},$
see Figure 7. Note that the family $\mathcal{H}$ depends on $h$, and the family $\mathcal{H}^{G}$ depends on $h$ and $\delta$; however, this dependency will be omitted to simplify the notation.
@@
with $C_{\Omega}$ a constant depending on $m_{d}(\Omega)$, where $m_{d}(\cdot)$ denotes the Lebesgue measure in $\mathbb{R}^{d}$. Taking into account that the number of edges of a $d$-dimensional hypercube is $2d(d-1)$, the Lebesgue measure $m_{d}(G^{h}_{\delta})$ of the grid $G_{\delta}^{h}$ intersecting $\mathcal{C}$ is bounded by
@@
Thus, for any $\gamma>0$, taking $\delta=h^{1+\gamma}$, the volume of $G_{\delta}^{h}$ tends to zero as $h\rightarrow 0$. In the following, we will take $\gamma=p$.
Figure 22. (A) Illustration of the simple function $f_{h}$ on $\mathcal{H}$, defined in (5.3). (B) The main features of $G^{h}_{\delta}$ are represented: The mesh thickness $\delta$, the size $h>0$, and the hyperrectangles $\mathcal{H}^{G}_{i}$, illustrated with orange boundary.
@@
That is, $f_{1,i}^{h}$ and $f_{2,i}^{h}$ are the average value of the function $f$ in the hyperrectangle $\mathcal{H}_{i}$ and $\mathcal{H}_{i}^{G}$, respectively. Then, we introduce the simple function
$f_{h}(x)=\sum_{i=1}^{N_{h}}f_{1,i}^{h}\chi_{\mathcal{H}_{i}}(x)+\sum_{i=1}^{N_{G}}f_{2,i}^{h}\chi_{\mathcal{H}_{i}^{G}}(x),$ (5.3)
where $\chi_{\mathcal{H}_{i}}$ denotes the characteristic function on the set $\mathcal{H}_{i}$, similar for $\mathcal{H}_{i}^{G}$. In the following, we denote by $M_{h}>0$ the number of values that the function $f_{h}$ takes on the family of hyperrectangles $\mathcal{H}$. Note that $M_{h}\leq N_{h}$.
@@
Moreover, as shown in [13, Section 6.2], if $f\in W^{1,p}(\Omega;\mathbb{R}_{+})$, there exists a constant $C>0$ independent of $h>0$ such that
@@
where $\text{diam}(\mathcal{H})=\max_{\mathcal{H}_{i}\in\mathcal{H}}\{\text{diam}(\mathcal{H}_{i})\}$. Since $\text{diam}(\mathcal{H^{G}})<\text{diam}(\mathcal{H})$, and $\text{diam}(\mathcal{H})$ is at most $\sqrt{d}h>0$, (5.5) reduces to
@@
$h_{1}=\frac{\varepsilon}{2C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}.$ (5.7)
@@
Here, we have one hyperplane for each canonical direction and one extra hyperplane in the direction $e_{\hat{\eta}}$, so there are in total $d+1$ hyperplanes. The constants $b_{k}$ are chosen in such a way that the hyperplanes are placed around the hyperrectangle $\mathcal{H}_{*}$, that is, $b_{k}$’s are taken such that
$\sigma(e_{\eta}x+b_{\eta})=0,\quad\sigma(-e_{\hat{\eta}}x+b_{\hat{\eta}}^{*})=0,\quad\text{for all }x\in\mathcal{H}_{*},\,\eta\in\llbracket 1,d\rrbracket.$
Let us apply the map defined by the parameters from the hyperplanes (5). This defines the new family of hyperrectangles $\{\mathcal{H}_{i}^{1}\}_{i=1}^{N}$. Namely, denote by $W^{1}\in\mathbb{R}^{d+1\times d}$ and $b^{1}\in\mathbb{R}^{d+1}$ the matrices given by
@@
Since (5.10), we have that either
$p_{\hat{\eta}}(x)=0\quad\text{or}\quad p_{d+1}(x)=0,\quad\text{for every }x\in\mathbb{R}^{d}.$ (5.13)
@@
We have that $F(\mathcal{H}_{*})={\bf 0}_{d}$, where ${\bf 0}_{d}$ denotes the null vector in $\mathbb{R}^{d}$. Moreover, for any hyperrectangles $\mathcal{H}_{i},\mathcal{H}_{j}\subset\mathcal{H}$ with $i\neq j$, we have that $F(\mathcal{H}_{i})\cap F(\mathcal{H}_{j})=\emptyset$.
@@
$=\left(\sum_{\eta=1}^{d}p_{\eta}(x)e_{\eta}x+p_{d+1}(x)e_{\hat{\eta}}x\right)+\left(\sum_{\eta=1}^{d}p_{\eta}(x)b_{\eta}e_{\eta}-p_{d+1}(x)b_{\hat{\eta}}^{*}e_{\hat{\eta}}\right)+b^{2}=F(x)+b^{2}.$
@@
Let $E$ represent the set of edges of $\mathcal{C}$. We define $\{E_{i}\}_{i=1}^{d}\subset E$ as a set of orthogonal edges. The family of hyperrectangles $\mathcal{H}^{E}$ is then given by $\mathcal{H}^{E}=\{\mathcal{H}_{i}\in\mathcal{H}\,:\,\mathcal{H}_{i}\cap\{E_{i}\}_{i=1}^{d}\neq\emptyset\},$ that is, the set of hyperrectangles in $\mathcal{H}$ intersecting at least one edge in $\{E_{i}\}_{i=1}^{d}$. We will prove that compressing the hyperrectangles in $\mathcal{H}^{E}$ into points also compresses all the hyperrectangles in $\mathcal{H}$ into points.
Figure 24. In the figure, we show an example of the compression process following the guidelines from Steps 2.1-2.2. First, in Step 1, we select two edges, $E_{1}$ and $E_{2}$ of $\mathcal{C}$, oriented according to different canonical vectors. Thus, $\mathcal{H}^{E}=\{\mathcal{H}_{1},\mathcal{H}_{4},\mathcal{H}_{7},\mathcal{H}_{8},\mathcal{H}_{9}\}$. Then, we compress the first hyperrectangle, $\mathcal{H}_{1}$, using two hyperplanes (a two-layer neural network with two neurons in the first layer and two neurons in the second layer). Here, $\mathcal{H}_{2}$ and $\mathcal{H}_{3}$ belong to $S_{\{1\}}$, and intersects $\mathcal{H}_{4}$ and $\mathcal{H}_{7}$ $S_{\{2\}}$. Next, in Step 2, we compress the second hyperrectangle on $E_{1}$, corresponding to $\mathcal{H}_{4}^{2}$. We use three hyperplanes to compress it (a two-layer neural network with three neurons in the first layer and two neurons in the second layer). Here, $\mathcal{H}_{5}^{2}$ and $\mathcal{H}_{6}^{2}$ belongs to $S_{\{1\}}$. In Step 3, we continue with $\mathcal{H}_{7}^{4}$. After compressing all the hyperrectangles along $E_{1}$, we have reduced every hyperrectangle to $(d-1)$-dimensional objects ($d-1=1$ in this example). By applying the same process to the hyperrectangles along $E_{2}$ and continuing with Steps 4 and 5, we complete the compression process.
@@
Let $E_{1}$ be one of the selected orthogonal edges of $\mathcal{C}$. By Step 2.1, we can compress any hyperrectangle $\mathcal{H}_{1}\in\mathcal{H}^{E}$ intersecting $E_{1}$ using a two-layer neural network. Moreover, observe that if a $d$-dimensional hyperrectangle $\mathcal{H}_{i}$ belongs to a subregion $S{I_{n}}$ with $n\leq d$, then it collapses into a $(d-1)$-dimensional hyperrectangle. Indeed, the condition $\mathcal{H}_{i}\subset S_{I_{n}}$ with $n\leq d$ implies that there exists an index $\bar{\eta}\in\llbracket 1,d+1\rrbracket$ such that $p_{\bar{\eta}}(x)=0$ for all $x\in\mathcal{H}_{i}$. Consequently, the $\bar{\eta}$-th coordinate of every point in $\mathcal{H}_{i}$ vanishes, collapsing $\mathcal{H}_{i}$ in to a $(d-1)-$dimensional hyperrectangle.
By applying the previous argument to all hyperrectangles in $\mathcal{H}^{E}$ intersecting $E_{1}$, we deduce that each $\mathcal{H}_{i}\in\mathcal{H}$ eventually enters a region $S_{I_{n}}$ with $n\leq d-1$. As a result, there exists a coordinate $\bar{\eta}\in\llbracket 1,d+1\rrbracket$ such that $p_{\bar{\eta}}(x)=0$ for all $x\in\mathcal{H}_{i}$, and thus $\mathcal{H}_{i}$ collapses into a $(d-1)$-dimensional hyperrectangle. Then, we continue with the hyperrectangles that intersect $E_{2}$. Since $E_{2}$ is orthogonal to $E_{1}$, the vanishing coordinate is now $\tilde{\eta}\neq\bar{\eta}$, and each $(d-1)$-dimensional hyperrectangle is further collapsed into a $(d-2)$-dimensional hyperrectangle. Repeating this process for each edge $E_{k}$ with $k\in\{1,\dots,d\}$, and noting that each step corresponds to the vanishing of a new and different coordinate, we conclude that every $\mathcal{H}_{i}\in\mathcal{H}$ is eventually mapped to a 00-dimensional object (a point). This concludes the compression process.
@@
If we had not followed the edge-based strategy, we would have needed to compress hyperrectangles located inside $\mathcal{C}$. In the worst case, enclosing a $d$-dimensional hyperrectangle requires $2d$ hyperplanes, implying that the neural network would need to have width $2d$. By restricting the compression to hyperrectangles on the edges of $\mathcal{C}$, we reduce this requirement to $d+1$ hyperplanes per step. Therefore, this strategy allows us to reduce the width of the neural network from $2d$ to $d+1$ neurons per layer.
@@
$\phi^{2N_{E}}(\mathcal{H}):=\phi(\mathcal{W}^{2N_{E}},\mathcal{B}^{2N_{E}},\mathcal{H})=\{x_{i}\}_{i=1}^{N_{h}},$ (5.17)
@@
$\phi^{L}(x_{i}):=\phi(\mathcal{W}^{L},\mathcal{B}^{L},x_{i})=f^{h}_{i},\quad\text{for all }i\in\llbracket 1,N_{h}\rrbracket.$ (5.18)
@@
$\phi^{\mathcal{L}}(x):=\phi(\mathcal{W}^{\mathcal{L}},\mathcal{B}^{\mathcal{L}},x)=f_{h}(x),\quad\text{for every }x\in\mathcal{H},$ (5.19)
@@
$\leq C_{1}^{p}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})\sum_{i=1}^{N_{h}^{G}}m_{d}(\mathcal{H}_{i}^{G})=C_{1}^{p}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})m_{d}(G^{h}_{\delta}).$
@@
with $C_{2}=C_{1}^{p}C_{\Omega,d}(1+\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}^{p})$. Then, since we are taking $\delta=h^{1+p}$, we obtain
@@
Let $\phi^{\mathcal{L}}$ be the map defined by (5.19), and denote by $l_{\mathcal{C}}$ the longest edge of $\mathcal{C}$. Then, for $h<l_{\mathcal{C}}\log(2)/(d+1)$ we have
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(\mathcal{C};\mathbb{R}_{+})}\leq C_{3}\left(1+\delta(h+\delta)+h\right),$
with $C_{3}>0$ a constant depending on $d$, $m_{d}(\mathcal{C})$, and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$.
Let us assume that $h<l_{\mathcal{C}}\log(2)/(d+1)$. Then, using Lemma 5.2, (5.2), and since $\delta<h^{\gamma+1}$ we deduce
$\|\phi^{\mathcal{L}}\|_{L^{\infty}(G_{\delta}^{h};\mathbb{R}_{+})}m_{d}(G^{h}_{\delta})$ $\leq C_{3}\left(1+\delta(h+\delta)+h\right)C_{\Omega,d}\delta(h+\delta)^{d-1}h^{-d}\leq 4C_{3}C_{\Omega,d}h^{p}.$ (5.23)
@@
$\|\phi^{\mathcal{L}}-f_{h}\|^{p}_{L^{p}(\Omega;\mathbb{R}_{+})}$ $\leq C_{2}h^{p}+(4C_{3}C_{\Omega,d}h^{p})^{p}\leq\hat{C}h^{p},$ (5.24)
where $\hat{C}=C_{2}+(4C_{3}C_{\Omega,d})^{p}$ is a positive constant depending on $d$, $p$, $m_{d}(\mathcal{C})$, and $\|f\|_{L^{p}(\Omega;\mathbb{R}_{+})}$. Thus, taking $h<\varepsilon/(2\hat{C}^{1/p})$, we deduce that
$\|\phi^{\mathcal{L}}-f_{h}\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\frac{\varepsilon}{2}.$ (5.25)
Therefore, taking $h<\min\{\varepsilon/(2\hat{C}^{1/p}),h_{1}\}$, with $h_{1}$ defined in (5.7), we have that (5.4) and (5.24) hold, and consequently, we conclude that
$\|f-\phi^{\mathcal{L}}\|_{L^{p}(\Omega;\mathbb{R}_{+})}<\varepsilon.$
Step 4 (Depth Estimation): To estimate the depth of the neural network, we first estimate $\min\{\varepsilon/(2\hat{C}^{1/p}),h_{1}\}$. Assume that $f\in W^{1,p}(\Omega;\mathbb{R}_{+})$. Then, let us observe that in an analogous way to (5.6), we can deduce that
@@
Therefore, using (5.26) instead (5.21), we have $C_{2}\leq C_{1}^{p}C_{\Omega,d}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$. Moreover, due to Lemma 5.2, we have that $C_{3}=C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$. Consequently, we deduce that the constant $\hat{C}$ can be taken as
$\hat{C}\leq(C_{1}^{p}C_{\Omega,d}+4CC_{\Omega,d})\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}=:\tilde{C}^{p}\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{p}$
Consequently, using (5.7) we have
$\min\left\{\frac{\varepsilon}{2\hat{C}^{1/p}},h_{1}\right\}$ $=\min\left\{\frac{\varepsilon}{2\hat{C}^{1/p}},\frac{\varepsilon}{2C\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}\right\}\geq\frac{\varepsilon}{2\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}\min\left\{\frac{1}{\tilde{C}},\frac{1}{C}\right\}$
$=:\frac{\varepsilon C_{5}}{2\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}.$
Then, in particular, we take
$h=\frac{\varepsilon C_{5}}{2\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}}.$ (5.27)
Now, due to Step 2.3, we know that $\mathcal{L}=2N_{h}+4M_{h}-1+2N_{E}$. Moreover, observe that $N_{E}$ and $N$ can be estimated by
@@
with $\delta=h^{1+p}$. Then, using $M_{h}\leq N_{h}$ and the estimations (5.28) and (5.27), we deduce the upper bound for the depth
$\mathcal{L}\leq C_{6}\left(\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}^{d}\varepsilon^{-d}+\|f\|_{W^{1,p}(\Omega;\mathbb{R}_{+})}\varepsilon^{-1}+1\right),$